# On the initial spin period distribution of neutron stars

### v1

@@ Appendix B Parametric functions for the spin period distribution
where $J$ and $K$ are known as the shape and scale parameters of the distribution, respectively. The subscript ‘$\nu$’ means that the corresponding distribution in frequency space is Weibull.
Appendix C Scale factor computations
@@
Here, $\rho_{\rm r}$ represents the space density for the galactocentric radius ($r$), and $\rho_{\rm z}$ is the density in the height above the Galactic plane ($z$). We assume that the pulsars distribute uniformly over the galactocentric azimuthal angle ($\phi$), which is justified since there is no apparent concentration of pulsars observed in the inner spiral arms of the Galaxy. The selection effects caused by the telescope’s limited sensitivity enter in the parameter $\eta(P,L,r,\phi,z)$ via the inverse-square law, $S_{\rm\nu}=Ld^{-2}$, where $S_{\rm\nu}$ is the apparent flux density detectable at a center frequency $\nu_{\rm center}$ (throughout we take $\nu_{\rm center}=1.4$ GHz). Given a minimum detectable flux density, $S_{\rm min}$, in a specific survey, $\eta(P,L,r,\phi,z)$ is set to 1 if $S_{\nu}\geq S_{\rm min}$ for a pulsar with period $P$ and luminosity $L$ at coordinates $\{r,\phi,z\}$, or otherwise it is set to 0. The numerator in Equation (C1) is integrated over the whole volume of the Galaxy, while the denominator is only integrated over the volumes where pulsars can be potentially detectable in any reference surveys.
The integral in Equation (C1) can be numerically computed at a number of values of $P$ utilizing Monte Carlo simulations. For each pulsar with $P$, $S_{\nu}$ and $S_{\rm min}$ are computed at a large number of grid points in the $\{r,\phi,z,L\}$ space. Then the numerator, which can be seen as the wighted total volume, is updated by adding the weight $\rho(r)\rho(z)$. The denominator, which can be thought to be the weighted sensitive volume, is updated by adding the weight $\rho(r)\rho(z)$ if $S_{\nu}\geq S_{\rm min}$. The settings are given below.
@@
$\rho(r)=37.6\left(\frac{r+0.55}{R_{\odot}+0.55}\right)^{1.64}\;\exp{\left[-4\left(\frac{r-R_{\odot}}{R_{\odot}+0.55}\right)\right]},$ (C2)
@@
$\rho(z)=0.75\exp{\left(-\frac{\lvert z\lvert}{0.18~{}\rm kpc}\right)}.$ (C3)
@@
$p(\log(L_{1400}))\propto\exp\left[-\frac{1}{2}\left(\frac{\log(L_{1400})-0.94}{0.85}\right)^{2}\right].$ (C4)
Figure 5: The distribution of the observed radio luminosities of 1834 pulsars with $P>10$ ms detected at 1400 MHz in the ATNF Catalogue. The red line is obtained by fitting the data to a log-normal distribution (see Equation C4).
@@
${\rm DM}=\int^{d}_{0}n_{e}{\rm d}l~{}\rm cm^{-3}\;pc,$ (C5)
[…]
@@
$S_{\rm min}(P)=\frac{\rho_{\rm s/n}\beta(T_{\rm r}+T_{\rm sky})}{G\sqrt{N_{\rm p}\Delta\nu t_{\rm int}}}\sqrt{\frac{W_{\rm e}}{P-W_{\rm e}}},\,P>W_{\rm e},$ (C6)
@@
$\frac{T_{\rm sky}(l,b,\nu)}{\rm K}=\left(25+\frac{275}{[1+(l/42)^{2}][1+(b/3)^{2}]}\right)\;\left(\frac{\nu}{408}\right)^{-2.6},$ (C7)
@@
$G_{\rm beam}(\mathcal{R})=G_{0}\;\exp\left(\frac{-2.77\mathcal{R}^{2}}{w^{2}}\right).$ (C8)
@@
$W_{\rm e}=\sqrt{W^{2}_{\rm int}+\tau^{2}_{\rm samp}+\tau^{2}_{\rm scatt}+\tau^{2}_{\rm DM}+\tau^{2}_{\rm\Delta DM}}.$ (C9)
@@
$\begin{split}\log\left(\frac{\tau_{\rm scatt}}{\rm ms}\right)=&-6.46+0.154\log{\rm DM}\\ &+1.07(\log{\rm DM})^{2}-3.86\log\left(\frac{\nu_{\rm center}}{\rm GHz}\right).\end{split}$ (C10)
@@
$\tau_{\rm DM}=8.3\rm\left(\frac{DM}{pc\;cm^{-3}}\right)\;\left(\frac{\Delta{\nu_{\rm chan}}}{MHz}\right)\;\left(\frac{\nu_{\rm center}}{GHz}\right)^{-3}~{}\mu s,$ (C11)
[…]
Appendix D Beaming fraction
@@
$f_{\rm TM98}(P>0.1~{}{\rm s})\approx 0.09~{}[\log\,(P/{\rm s})-1]^{2}+0.03.$ (D2)
[…]
$f_{\rm TM98}(P\leq 0.1~{}{\rm s})\approx-0.255~{}\log(P/{\rm s})+0.135.$ (D3)

### v2

@@ Appendix B Parametric functions for the spin period distribution
where $J$ and $K$ are known as the shape and scale parameters of the distribution, respectively. The subscript ‘$\nu$’ means that the corresponding distribution in frequency space is Weibull. The Weibull distribution peaks at $K^{-1}[J/(J+1)]^{1/J}$.
Appendix C Comparison with previous works
Figure 6 displays the initial periods data along with the log-normal distributions obtained by Igoshev et al. (2022) (in red) and inferred in this work (in green); both without the correction of selection effects. One can see that both sets of results are generally consistent with each other. However, in our work the log-normal model is disfavored. This is due to a different analysis method, and more importantly different data selection criteria, especially the treatment of the observed pulsars with supposedly imaginary initial periods (see Section 2.2).
Figure 6: The log-normal distribution obtained in this work (green) and that in Igoshev et al. (2022) (red). The histograms show the initial periods used by Igoshev et al. (2022) (in blue) along with that analyzed in this work (in orange).
Appendix D Scale factor computations
@@
Here, $\rho_{\rm r}$ represents the space density for the galactocentric radius ($r$), and $\rho_{\rm z}$ is the density in the height above the Galactic plane ($z$). We assume that the pulsars distribute uniformly over the galactocentric azimuthal angle ($\phi$), which is justified since there is no apparent concentration of pulsars observed in the inner spiral arms of the Galaxy. The selection effects caused by the telescope’s limited sensitivity enter in the parameter $\eta(P,L,r,\phi,z)$ via the inverse-square law, $S_{\rm\nu}=Ld^{-2}$, where $S_{\rm\nu}$ is the apparent flux density detectable at a center frequency $\nu_{\rm center}$ (throughout we take $\nu_{\rm center}=1.4$ GHz). Given a minimum detectable flux density, $S_{\rm min}$, in a specific survey, $\eta(P,L,r,\phi,z)$ is set to 1 if $S_{\nu}\geq S_{\rm min}$ for a pulsar with period $P$ and luminosity $L$ at coordinates $\{r,\phi,z\}$, or otherwise it is set to 0. The numerator in Equation (D1) is integrated over the whole volume of the Galaxy, while the denominator is only integrated over the volumes where pulsars can be potentially detectable in any reference surveys.
The integral in Equation (D1) can be numerically computed at a number of values of $P$ utilizing Monte Carlo simulations. For each pulsar with $P$, $S_{\nu}$ and $S_{\rm min}$ are computed at a large number of grid points in the $\{r,\phi,z,L\}$ space. Then the numerator, which can be seen as the wighted total volume, is updated by adding the weight $\rho(r)\rho(z)$. The denominator, which can be thought to be the weighted sensitive volume, is updated by adding the weight $\rho(r)\rho(z)$ if $S_{\nu}\geq S_{\rm min}$. The settings are given below.
@@
$\rho(r)=37.6\left(\frac{r+0.55}{R_{\odot}+0.55}\right)^{1.64}\;\exp{\left[-4\left(\frac{r-R_{\odot}}{R_{\odot}+0.55}\right)\right]},$ (D2)
@@
$\rho(z)=0.75\exp{\left(-\frac{\lvert z\lvert}{0.18~{}\rm kpc}\right)}.$ (D3)
@@
$p(\log(L_{1400}))\propto\exp\left[-\frac{1}{2}\left(\frac{\log(L_{1400})-0.94}{0.85}\right)^{2}\right].$ (D4)
Figure 7: The distribution of the observed radio luminosities of 1834 pulsars with $P>10$ ms detected at 1400 MHz in the ATNF Catalogue. The red line is obtained by fitting the data to a log-normal distribution (see Equation D4).
@@
${\rm DM}=\int^{d}_{0}n_{e}{\rm d}l~{}\rm cm^{-3}\;pc,$ (D5)
where $d$ is in units of pc and $n_{e}$ denotes the free electron density in cm-3. To account for the errors in estimating the model dispersion measure (DMmod) by using this electron density model, we take a random error drawn from a Gaussian distribution with a mean of zero and width of $0.2\,\rm DM_{\rm mod}$.
@@
$S_{\rm min}(P)=\frac{\rho_{\rm s/n}\beta(T_{\rm r}+T_{\rm sky})}{G\sqrt{N_{\rm p}\Delta\nu t_{\rm int}}}\sqrt{\frac{W_{\rm e}}{P-W_{\rm e}}},\,P>W_{\rm e},$ (D6)
@@
$\frac{T_{\rm sky}(l,b,\nu)}{\rm K}=\left(25+\frac{275}{[1+(l/42)^{2}][1+(b/3)^{2}]}\right)\;\left(\frac{\nu}{408}\right)^{-2.6},$ (D7)
@@
$G_{\rm beam}(\mathcal{R})=G_{0}\;\exp\left(\frac{-2.77\mathcal{R}^{2}}{w^{2}}\right).$ (D8)
@@
$W_{\rm e}=\sqrt{W^{2}_{\rm int}+\tau^{2}_{\rm samp}+\tau^{2}_{\rm scatt}+\tau^{2}_{\rm DM}+\tau^{2}_{\rm\Delta DM}}.$ (D9)
@@
$\begin{split}\log\left(\frac{\tau_{\rm scatt}}{\rm ms}\right)=&-6.46+0.154\log{\rm DM}\\ &+1.07(\log{\rm DM})^{2}-3.86\log\left(\frac{\nu_{\rm center}}{\rm GHz}\right).\end{split}$ (D10)
@@
$\tau_{\rm DM}=8.3\rm\left(\frac{DM}{pc\;cm^{-3}}\right)\;\left(\frac{\Delta{\nu_{\rm chan}}}{MHz}\right)\;\left(\frac{\nu_{\rm center}}{GHz}\right)^{-3}~{}\mu s,$ (D11)
where the channel bandwidth $\Delta{\nu_{\rm chan}}\ll\nu_{\rm center}$ is assumed; for the 1.4 GHz surveys, we employ $\Delta{\nu_{\rm chan}}=100/256$ MHz for $\rm DM<100$ cm-3 pc and $\Delta{\nu_{\rm chan}}=100/512$ MHz for $\rm DM>100$ cm-3 pc. The last term, $\tau_{\Delta\rm DM}$, is the smearing due to the finite DM step size in the survey, which is neglected here.
Pulsar survey. We use the sensitivity limit of the Parkes Multibeam Survey (Manchester et al., 2001) to evaluate the scale factors, its parameters are shown in Table 5. The panel (a) of Figure 4 displays the sensitivity curve of this survey. Note that, to obtain the $S_{\rm min}$ in this plot, we have simply adopted a sky temperature of 25 K, the observed pulse width is approximated as $[(0.04P)^{2}+t^{2}_{\rm samp}+\tau^{2}_{\rm DM}]^{1/2}$, and $\tau_{\rm DM}\approx t_{\rm samp}\frac{\rm DM}{\rm DM_{0}}$ with $\rm DM_{0}=28$ cm-3 pc.
Appendix E Beaming fraction
@@
$f_{\rm TM98}(P>0.1~{}{\rm s})\approx 0.09~{}[\log\,(P/{\rm s})-1]^{2}+0.03.$ (E2)
While this relation was derived based on the radio pulsars with luminosities $>0.1$ mJy kpc2 and might differ for dimmer objects, we assume it is valid for the dimmer pulsars. We extend Equation (E2) linearly to estimate the beaming fraction for pulsars with periods $P\leq 0.1$ s: with a lower limit of $f_{\rm TM98}(P=0.1~{}{\rm s})\approx 0.39$ and the upper limit of 0.9 obtained from millisecond pulsars (Kramer et al., 1998), this relation can be extended to the formula (Titus et al., 2020)
$f_{\rm TM98}(P\leq 0.1~{}{\rm s})\approx-0.255~{}\log(P/{\rm s})+0.135.$ (E3)